Stellar Lumber Company handles three principal lines of merchandise with these varying rates of gross profit on cost.

Lumber 25%
Millwork 30%
Hardware and fittings 40%


On August 18, a fire destroyed the office, lumber shed, and a considerable portion of the lumber stacked in the yard. To file a report of loss for insurance purposes, the company must know what the inventories were immediately preceding the fire. No detail or perpetual inventory records of any kind were maintained. The only pertinent information you are able to obtain are the following facts from the general ledger, which was kept in a fireproof vault and thus escaped destruction.

Lumber

Millwork

Hardware

Inventory, Jan. 1, 2020 $249,800 $88,600 $45,800
Purchases to Aug. 18, 2020 1,488,200 380,300 162,900
Sales to Aug. 18, 2020 2,055,600 494,000 211,400


Submit your estimate of the inventory amounts immediately preceding the fire.
Answer #1

Particulars

lumber millwork hardware
opening stock $249800 $88600 $45800
Add purchase $1488200 $380300 $162900
Less cost of goods sold ($1644480) ($380000) ($151000)
Closing stock $93520 $88900 $57700

Note calculation of cost of goods sold

Lumber = $2055600*100/125 = $1644480

Millwork = $494000*100/130 = $380000

Hardware = $211400*100/140 = $151000

Feel free to ask any queries..

Also plz upvote.. it means a lot..